Find the equation of Line passing through the points (a cos x, a sin x) and (a sin y, a sin y)

Grade:11

1 Answers

Arun
6 years ago
Dear student
Equation of line is given as
y - y1 = m (x- x1)
Where
m = y2 - y1/(x2 - x1)
Hence equation of line:
y - a sin x = [(a siny - a sinx)/( a siny - a cosx)]   (x - a cosx)
Now simplify it.
Regards
Arun (askIITians forum expert)